The vendor provided 95% pure antimony trioxide [Sb2O3] powder chemical. Molar mass of elements (Sb: 121.8 g/mol; O: 16 g/mol). Solve the following problems to get accurate initial Sb(III) concentration for lab-scale experiments. Hint: The Sb(III) solution needs to be prepared in 1 M HCl solution.

i.How much chemical in mg/L will be required for preparation of 1 g/L Sb(III) stock solution?

ii.What will be the dilution factor to prepare 10 mg/L Sb(III) solution?


1
Expert's answer
2022-02-28T11:09:46-0500

95% pure antimony trioxide [Sb2O3]


Sb: 121.8 g/mol


O: 16 g/mol


1 M HCl solution


(i.)Chemical in mg/L will be required for preparation of 1 g/L Sb(III) stock solution


=121.8g/mol+16 g/mol=137.8 g/mol


=137.8 g/mol"\times" 0.95


=130.91mg/L


(ii.) The dilution factor to prepare 10 mg/L Sb(III) solution


="\frac{130.91mg/L \times× 0.05}{0.95}"


= 6.89